LeRoy has 15 different ties. He chooses a tie at random everyday.
How many times could he expect to wear a given tie in 45 days?

Answer by stanbon(75887)

LeRoy has 15 different ties. He chooses a tie at random everyday.
How many times could he expect to wear a given tie in 45 days?
---
Binomial Problem with n = 45 and p = 1/15
---
Expected value = np = 45*(1/15) = 3 times
============================================
Cheers,
Stan H.